OpenStudy (anonymous):

I Dont Understand Write the equation of the line that passes through (-4, 12) and (7, 23). A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (anonymous):

go to math way its a free site and its amazing it gives you answers and plots points on graphs and everything

OpenStudy (anonymous):

First we have to use this formula to find the m\[\color{green}{y}-\color{red}{y_o}=m*(\color{green}{x}-\color{red}{x_o})\]Where\[(\color{green}{x,y})=(-4,12)\]and\[(\color{red}{x_o,y_o})=(7,23)\]Replacing, we've got:\[\boxed{m=1}\]Now you can chose one point to make the equation of the line, I'll chose the first one\[(\color{red}{x_o,y_o})=(-4,12)\]\[y-12=1*(x-(-4))\]\[y=x+4+12\]The answer is:\[\boxed{y=x+16}\]

OpenStudy (anonymous):

its C